# Natural Language to Verilog: Design of a Recurrent Spiking Neural Network using Large Language Models and ChatGPT

### v2

@@ Abstract
This paper investigates the use of Large Language Models (LLMs) and natural language prompts to generate hardware description code, namely Verilog. Building on our prior work, we employ OpenAI’s ChatGPT4 and natural language prompts to synthesize an RTL Verilog module of a programmable recurrent spiking neural network, while also generating test benches to assess the system’s correctness. The resultant design was validated in three simple machine learning tasks, the exclusive OR, the IRIS flower classification and the MNIST hand-written digit classification. Furthermore, the design validated on on a Field-Programmable Gate Array (FPGA) and subsequently synthesized in the SkyWater 130 nm technology by using an open-source electronic design automation flow. The design was submitted to Efabless Tiny Tapeout 6.
@@ I Introduction
To address these challenges, generative Artificial Intelligence (GenAI) and specifically Large Language Models (LLMs), could augment state-of-the-art Electronic Design Automation (EDA) tools, and offer viable solutions [8]. Specifically, in this paper we demonstrate that it is possible to use natural language prompts to generate hardware description code, namely Verilog. Since the advent of LLMs in the 4Q of 2023, there has been efforts to use LLMs in the hardware design to streamline and accelerate the process [9, 10]. Tailoring for hardware debugging, Fu et al. used them for the identification and resolutions of bugs in hardware designs [11, 12]. The authors in [13, 14, 15] employ “aligned” LLMs to generate hardware description code in Verilog or VHDL with reduced syntax errors. The most popular choice for LLM based hardware design has been OpenAI’s GPT-4. J. T. Meech in [16] used Microsoft Bing-chat, which uses GPT-4, in order to generate a permuted congruential random number generator design with a wishbone interface. The die-harder test and the simulations he conducted verified its functionality and quality of the design. Similarly, our previous work tackled a programmable 3 by 3 digital spiking neuron array created from conversational language using GPT-4. The above implementation was submitted for fabrication in Skywater 130 nm technology and was verified through simulation [17].
@@
In this paper we build on our prior work [17], and explore GPT-4’s potential to design a more complex hardware design architecture, namely a 3 layer, 3 neuron per layer Recurrent Spiking Neuron Network (RSNN) w. Furthermore we demonstrate how to use natural language prompts and LLMs to support test-bed development and hence the verification process.

### v3

@@ Abstract
This paper investigates the use of Large Language Models (LLMs) and natural language prompts to generate hardware description code, namely Verilog. Building on our prior work, we employ OpenAI’s ChatGPT4 and natural language prompts to synthesize an RTL Verilog module of a programmable recurrent spiking neural network, while also generating test benches to assess the system’s correctness. The resultant design was validated in three simple machine learning tasks, the exclusive OR, the IRIS flower classification and the MNIST hand-written digit classification. Furthermore, the design was validated on a Field-Programmable Gate Array (FPGA) and subsequently synthesized in the SkyWater 130 nm technology by using an open-source electronic design automation flow. The design was submitted to Efabless Tiny Tapeout 6.
@@ I Introduction
To address these challenges, generative Artificial Intelligence (GenAI) and specifically Large Language Models (LLMs), could augment state-of-the-art Electronic Design Automation (EDA) tools, and offer viable solutions [8]. Specifically, in this paper we demonstrate that it is possible to use natural language prompts to generate hardware description code, namely Verilog. Since the advent of LLMs in the 4Q of 2023, there has been efforts to use LLMs in the hardware design to streamline and accelerate the process [9, 10, 11, 12]. J. T. Meech in [13] used Microsoft Bing-chat in order to generate a permuted congruential random number generator design with a wishbone interface. Tailoring for hardware debugging, Fu et al. used them for the identification and resolutions of bugs in hardware designs [14]. The most popular choice for LLM based hardware design has been OpenAI’s GPT. The authors in [15, 16] employed GPT to generate hardware description code in Verilog or VHDL, resulting in a reduction of syntax errors. Similarly, our previous work tackled a programmable 3 by 3 digital spiking neuron array created from conversational language using GPT-4. The above implementation was submitted for fabrication in Skywater 130 nm technology and was verified through simulation [17].
@@
In this paper we build on our prior work [17], and explore GPT-4’s potential to design a more complex hardware design architecture, namely a 3 layer, 3 neuron per layer Recurrent Spiking Neuron Network (RSNN). Furthermore we demonstrate how to use natural language prompts and LLMs to support test-bed development and hence the verification process.